OpenStudy (anonymous):

How many integers are between 99 and 201 are divisible by 2 or 5

OpenStudy (anonymous):

first no is 100 and last no. is 200 1. a=100,d=2,l=200 find n1 by l=a+(n1-1)d,find n1 2. a=100,d=5,l=200 find n2 by l=a+(n2-2)d, 3. L.C.M of 2&5 is 10 a=100,d=10,l=200 find n3 by l=a+(n3-1) d Required no.is=n1+n2-n3

OpenStudy (anonymous):

i solve for n1 200=100+(n1-1)*2 200-100=(n1-1)*2 100/2=n1-1 50+1=n1 n1=51 similarly you can solve for n2 and n3.